# NavGPT-2: Unleashing Navigational Reasoning Capability for Large Vision-Language Models

### v1

@@ 2 Related Works
Vision-and-Language Navigation (VLN) The pursuit of developing a universal navigation agent capable of following free-form linguistic directives to navigate within an unfamiliar photorealistic environment has been a longstanding objective in the field of Vision-and-Language Navigation [7, 33, 56, 65]. These methods solve the task from two main aspects. (1) Vision-Language alignment. Some work [27, 28, 26, 51, 13, 15, 57, 39] benefits from generic visual-linguistic representations [17, 40, 62, 37, 63], some exploit additional supervision from data augmentation [4, 64, 35, 54, 75, 26, 22, 69, 34], training stragies [73, 72, 49, 83, 29] to learn such cross-modality alignment. (2) Efficient action planning mechanism with historical state memorization [28, 14], self-correction [32, 50], navigation map construction [2, 12, 16, 70, 79, 44, 74], and external knowledge prompts [42, 38]. In this work, we investigate the method of boosting a simplified prevalent VLN policy with visual-linguistic representations in LLM’s latent space.
@@ 4.2 Comparison with State of the Art
As shown in Table 1, we list our model with a similar size222Our model is smaller (1.5B and 5B) than original FlanT5 models (3B and 11B) as we only utilize the LLM encoder during navigation. as previous LLM-based methods, (1.5B) and (5B). Our 5B model outperforms NaviLLM [80] (7B) by 3% SR on test split while still maintaining the language capacity to generate self-explained navigation reasoning. NavGPT-2 far exceeding NavGPT [81] and surpasses all the zero-shot methods, since those methods rely on extensive prompt engineering and struggle to model navigation history. Compared to the baseline methods, NavGPT-2 bypass it by 4% SR and 2% SPL on the test split even if we do not incorporate with VLN pertaining. The current SOTA method [75] is achieved by scaling up the training environment for DUET with HM3D [60] and Gibson [76], besides the original 61 scenes in MP3D [9]. When considering the same training scale in MP3D, our method beats the original DUET by in SR on the test unseen split and is comparable with metric map-based method [3] which incorporates an extra depth sensor to construct Bird’s-Eye-View (BEV) perception.

### v2

@@ 2 Related Works
Vision-and-Language Navigation (VLN) The pursuit of developing a universal navigation agent capable of following free-form linguistic directives to navigate within an unfamiliar photorealistic environment has been a longstanding objective in the field of Vision-and-Language Navigation [7, 33, 56, 65]. These methods solve the task from two main aspects. (1) Vision-Language alignment. Some work [27, 28, 26, 51, 13, 15, 57, 39, 79] benefits from generic visual-linguistic representations [17, 40, 62, 37, 63], some exploit additional supervision from data augmentation [4, 64, 35, 54, 75, 26, 22, 69, 34], training stragies [73, 72, 49, 84, 29] to learn such cross-modality alignment. (2) Efficient action planning mechanism with historical state memorization [28, 14], self-correction [32, 50, 79], navigation map construction [2, 12, 16, 70, 80, 44, 74], and external knowledge prompts [42, 38]. In this work, we investigate the method of boosting a simplified prevalent VLN policy with visual-linguistic representations in LLM’s latent space.
@@ 4.2 Comparison with State of the Art
As shown in Table 1, we list our model with a similar size222Our model is smaller (1.5B and 5B) than original FlanT5 models (3B and 11B) as we only utilize the LLM encoder during navigation. as previous LLM-based methods, (1.5B) and (5B). Our 5B model outperforms NaviLLM [81] (7B) by 4% SR on test split while still maintaining the language capacity to generate self-explained navigation reasoning. NavGPT-2 far exceeding NavGPT [82] and surpasses all the zero-shot methods, since those methods rely on extensive prompt engineering and struggle to model navigation history. Compared to the baseline methods, NavGPT-2 bypass it by 5% SR and 2% SPL on the test split even if we do not incorporate with VLN pertaining. The current SOTA method [75] is achieved by scaling up the training environment for DUET with HM3D [60] and Gibson [76], besides the original 61 scenes in MP3D [9]. When considering the same training scale in MP3D, our method beats the original DUET by in SR on the test unseen split and is comparable with metric map-based method [3] which incorporates an extra depth sensor to construct Bird’s-Eye-View (BEV) perception.